### v2

@@ Introduction
To help to address this lack of 3D training data, Ashton et al. [ashton2024ahmed] recently created the AhmedML dataset, that contains 500 different geometry variants of the Ahmed car body, simulated using a high-fidelity hybrid RANS-LES approach using OpenFOAM on meshes of approximately 20M cells. This open-source dataset contains time-averaged volume & boundary flow-field variables as well forces and moment data to provide a rich dataset to aid ML model development. Whilst this dataset enables the development and testing of ML methods, there are a number of limitations to this dataset that have motivated the creation of the WindsorML dataset.
@@
Finally, the choice of the Windsor body is also motivated by its use within the Automotive CFD Prediction Workshops (AutoCFD) 111https://autocfd.org that aim to bring the fluid dynamics and automotive community together to improve the state of the art for CFD prediction of road vehicles. The dataset described in this paper was used as a test-case during the 4th AutoCFD workshop, for which future papers will discuss the resulting comparisons from a range of different ML methods. Concurrently the DrivAerML dataset [ashton2024drivaer] based upon the open-source Drivaer has been created, that is described in it’s own paper. Used together, the AhmedML [ashton2024ahmed], WindsorML and DrivAerML [ashton2024drivaer] datasets, provide ML developers with a broad set of data to use that has been formatted in a consistent fashion and is openly available to download and use with a permissive license (CC-BY-SA).
@@
The Windsor body was created by Steve Windsor and Jeff Howell of Rover and later Jaguar Land Rover (JLR) [Varney2020a, page2022] to improve upon the well studied Ahmed car body [Lienhart2003, Ahmed1984, meile2011, Ashton2015a] by creating a geometry whose flow features more closely a road vehicle. Whilst the Ahmed car body has been used in many experimental and CFD studies, the long flat surface between the front and rear of the body decouples the aerodynamic behaviour of these two regions and makes it more relevant to commercial vehicles. Two variants of the Windsor body were experimentally tested at the Loughborough University wind-tunnel [Varney2020a] and have been the subject of numerous CFD studies - that most recently have taken place within the 2nd, 3rd and 4th Automotive CFD Prediction Workshops [page2022]. The Windsor body was tested in numerous configurations but for this study we follow the 4th AutoCFD4 setup, where the vehicle has no wheels (which simplifies a common source of uncertainty) and the squareback rear shape is used, as the baseline geometry. As discussed later, this baseline geometry is then adapted according to 7 parameters to generate 355 different geometries.
@@
The model geometry is shown in Figure 1. The reference frontal area is defined by the vehicle height and width and for the baseline geometry is . The reference length used for pitching moment is the wheelbase . The CAD geometry of the model has its origin on the ground plane, in the symmetry plane midway between the wheels (which are not included in this particular geometry) i.e . This is also the moment reference center for the calculating of the moments.
@@
All simulations were run using the Volcano ScaLES code by Volcano Platforms, which solves the explicit compressible Navier–Stokes equations using a nominally 4th order spatially accurate finite difference discretization with favorable Kinetic Energy and Entropy Consistency properties making it suitable for Large Eddy Simulations (LES) over a range of flow regimes. The viscous flux discretization utilizes a mix of 4th order and 2nd order discretizations with high spectral bandwidth thereby allowing further robustness for high Reynolds number flows without additional numerical dissipation via the inviscid flux numerical operator in turbulence resolving regions. Physics-based numerical sensors that are functions of local velocity gradients as well as pressure and density fluctuations allow for spatio-temporally localized use of limiters needed to capture flows with discontinuities such as shocks and steep density gradients. The Strong Stability Preserving (SSP) variant of the classical 3rd order Runge–Kutta scheme developed by Gottlieb & Shu[gottlieb1998total] is utilized for all the work presented in this paper.
Geometries are represented in the numerical formulation using an immersed boundary algorithm capable of enforcing inviscid no-penetration boundary conditions as well as viscous skin friction with appropriate Reynolds number asymptotic properties using an equilibrium wall-model. The wall-model that provides a shear-stress constraint at the wall uses solution information interpolated via probing at a fixed distance of into the fluid away from walls. Additional methods to damp generation of spurious numerical noise at walls as well as at coarse-fine interfaces are also utilized in the present work. All simulations presented in this work utilize the constant coefficient Vreman subgrid scale closure[vreman2004eddy] to model the subgrid-scale stresses.
@@
To follow the setup of the AutoCFD 3 and 4 workshops, the Windsor body is yawed 2.5 degrees and the Reynolds number is based on the body length and the freestream velocity. A nominal inlet velocity of is given upstream of the car and a sensor is placed at to measure local static pressure and velocity. For the validation of the method against experimental data for the baseline geometry, these are used to normalize the simulation outputs, however for the dataset itself, the reference inlet velocity is used to normalize all outputs (see SI for further details). The remaining boundary conditions are shown in Figure 2.
@@
Only the ground plane has a no slip condition and hence has boundary layer growth; the top and side walls should be treated as a slip or ‘inviscid’ wall.
@@
All simulations were run on Amazon Web Services, using a dynamic HPC cluster provisioned by AWS ParallelCluster v3.9. Amazon EC2 g5.48xlarge nodes were used for each case, which contain 8 Nvidia A10g GPU cards and a second generation AMD EPYC processor with 96 physical cores and 768GB of CPU memory. These are connected using a 100Gbit/s Elastic Fabric Adapter (EFA) interconnect [Shalev2020]. A 300TB Amazon Fsx for Lustre parallel file-system was used as the main data location during the runs, which were later transferred to object storage in Amazon S3. Each simulation was run on a single g5.48xlarge node i.e 8 A10g GPUs, and each case took approximately 2mins to mesh and 28hrs to solve.
@@
| Label | Min. | No. of Cells | Wall time / CTU | |
@@
For the design case study, we choose to keep the refinement regions fixed for all design cases. Using the fixed refinement topology, we consider a consistent grid refinement study for the Windsor body. An example of a grid in this family is depicted in Figure 3. The surface of the body is refined to the finest level with a large wake block immediately behind the body. The block is taller than necessary for the Windsor body but since we are choosing to keep the refinement regions fixed we define a region to accommodate all designs ride height. Table 1 shows results of a consistent grid refinement study using a sequence of grids for and . The columns of the table show the minimum grid spacing used, the number of computational cells in the grid, the wall-clock time per convective time unit222Nondimensional time is defined to be where is the length of the Windsor body. (CTU) in minutes and the drag coefficient . An additional row is included for the experimental drag coefficient value. Based on the results, is chosen as a representative baseline case and the remainder of the experimental comparisons are carried out for the grid. This was to achieve a balance between accuracy, computational cost and reach a mesh size that is typically used by the automotive industry for external aerodynamics.
Velocity: Typical of bluff-bodies, the Windsor body shows a large recirculation region at the base. The data for the slice was taken with a freestream velocity of which is the same as our simulation but, in addition to this plane, tomographic PIV was taken for a 3D region behind the car. This data was recorded for a flow with a freestream velocity of . To make comparisons with the tomographic PIV data, we normalize the velocities by their freestream velocity. With that caveat, Figures 4 & 5 shows a comparison of the normalized time-averaged streamwise velocity for compared with the PIV measurements and the same solution at compared to the tomographic PIV data. The overall shape of the wake region is well captured as well as its interaction with the floor boundary layer which can be seen in the velocity contours. See SI for a detailed discussion on the validation of the underlying CFD method.
@@
The dataset is open-source under the CC-BY-SA license 333https://creativecommons.org/licenses/by-sa/4.0/) and available to download via Amazon S3. In order to download, no AWS account is required and the full details are provided in the SI and the dataset README 444https://caemldatasets.s3.us-east-1.amazonaws.com/windsor/dataset/README.txt.
@@
The dataset follows the same structure as two concurrently developed datasets; AhmedML [ashton2024ahmed] and DrivAerML [ashton2024drivaer]. A summary of the dataset is provided in Table 3 (see SI for full details). The purpose of the dataset is to provide a rich dataset to support the development of a broad range of potential ML approaches. For this reason, we provide all possible outputs such as the full volume flow-field, the boundary surface, images of the flow-field as well as the time-averaged force and moment coefficients. All the outputs are provided in commonly used open-source formats (vtp, vtu etc) which allow for the the broadest compatibility and ease of use for developers.
@@ A.1 Large Eddy Simulations
Note that the dynamic viscosity, can be expressed as a function of Favre-filtered static temperature using an appropriate law (such as the Sutherland’s law). However, in the present work, this dynamic viscosity has be assumed to a spatio-temperal constant with a value determined to achieve the required Reynolds number. Static temperature, is a function of the local pressure and density in accordance to the ideal gas law using a gas constant, :
@@
where the unclosed terms on the right hand side represent interactions between subgrid-scale energy and the resolved scale energy. Since these terms are only relevant at high Mach numbers, the residual terms in the energy equation are modelled via just a turbulent Prandtl number and the Vreman’s eddy viscosity model used for the momentum equations. The positive definiteness of the SGS eddy viscosity ensures that the numerical model for the filtered energy equation is purely dissipative, i.e. it assumes that the energy is exclusively transfered from the large scales to the small scales of motion.
@@ B.1 Licensing terms
The dataset is provided with the Creative Commons CC-BY-SA v4.0 license555https://creativecommons.org/licenses/by-sa/4.0/deed.en. A full description of the license terms is provided under the following URL:
@@ B.3 Long-term hosting/maintenance plan
The data is hosted on Amazon S3 as it provides high durability, fast connectivity and is accessible without first requesting an account or credentials (only AWS CLI tools, described above, which are free to download and use). A dedicated website is currently being created for this dataset and the two associated datasets; AhmedML [ashton2024ahmed] and DrivAerML datasets [ashton2024drivaer] to help further clarify where the data is hosted and to communicate any additional mirroring sites.
@@ B.7 Geometry variants
The geometry choice within the dataset results in a broad range of flow physics - that is partly illustrated in Figures 11 and 12 which show the large range of drag and lift coefficients across the dataset. Figures 9 and 10 take two examples from the dataset that shows the resulting flow-field (mean streamwise velocity) for a geometry producing a high drag coefficient and another one for a low drag coefficient.
@@
Figures 13, 14, 15, & 16 are provided to give an illustration of the range of flow features and post-processing that makes this a rich dataset for the development and testing of ML methods.
@@ Appendix C Additional Validation details
Pressure: In Figure 17 and Figure 18, we show experimental and computed surface pressure coefficient values, , for the cut planes and respectively. For the plane we observe stronger suction in the CFD compared to the experiment near where the flow transitions, which is typical in other CFD simulations for this problem. Otherwise, there is good agreement for the upper and lower sections. For the cut, the CFD again shows a stronger suction peak this time in the A-pillar area. Small variations in the data suggest a longer averaging could be done but it would not be expected to change the agreement. Overall, the agreement is reasonably good and similar to other results seen in the AutoCFD3 results. A comparative view of the base pressure is shown in Figure 19 where similar pressure distributions are seen with a low pressure region on the leeward side of the vehicle associated with the recirculation region. The simulation results show a higher pressure that the experiment which is most evident on the side, however the spatial distribution of the pressure and thus the shape of the contours match well.
@@
Velocity: Time-averaged streamlines colored by normalized mean streamwise velocity at are shown in Figure 20 compared with experimental data from the thesis by Varney [Varney2020]. The streamlines show the large recirculation region behind the vehicle and the correct flow topology. The data for the slice was taken with a freestream velocity of which is the same as our nominal freestream velocity but, in addition to this plane, tomographic PIV was taken for a 3D region behind the car. This data was recorded for a flow with a freestream velocity of . To make comparisons with the tomographic PIV data, we normalize the velocities by their freestream velocity. Figures 21 and 22 show progressive cuts through the wake going downstream. Starting nearest the car, the experiment shows a thicker shear layer but as we move downstream the predicted and experimental flow have essentially the same shear layer thickness. The overall shape of the wake region is well captured as well as its interaction with the floor boundary layer which can be seen in the velocity contours.
Appendix D Datasheet
D.1 Motivation

### v3

@@ Introduction
To help to address this lack of 3D training data, Ashton et al. [ashton2024ahmed] recently created the AhmedML dataset, that contains 500 different geometry variants of the Ahmed car body, simulated using a high-fidelity hybrid RANS-LES approach using OpenFOAM on meshes of approximately 20M cells. This open-source dataset contains time-averaged volume & boundary flow-field variables as well as forces and moment data to provide a rich dataset to aid ML model development. Whilst this dataset enables the development and testing of ML methods, there are a number of limitations to this dataset that have motivated the creation of the WindsorML dataset.
@@
Finally, the choice of the Windsor body is also motivated by its use within the Automotive CFD Prediction Workshops (AutoCFD) 111https://autocfd.org that aim to bring the fluid dynamics and automotive community together to improve the state of the art for CFD prediction of road vehicles. The dataset described in this paper was used as a test-case during the 4th AutoCFD workshop, for which future papers will discuss the resulting comparisons from a range of different ML methods. Concurrently the DrivAerML dataset [ashton2024drivaer] based upon the open-source Drivaer has been created, that is described in it’s own paper. Used together, the AhmedML [ashton2024ahmed], WindsorML and DrivAerML [ashton2024drivaer] datasets, provide ML developers with a broad set of data to use that has been formatted in a consistent fashion and is openly available to download and use with a permissive license (CC-BY-SA). A dedicated website 222https://caemldatasets.org for these datasets has been created to provide the community with up to date information.
@@
The Windsor body was created by Steve Windsor and Jeff Howell of Rover and later Jaguar Land Rover (JLR) [Varney2020a, page2022] to improve upon the well studied Ahmed car body [Lienhart2003, Ahmed1984, meile2011, Ashton2015a] by creating a geometry whose flow features more closely match a road vehicle. Whilst the Ahmed car body has been used in many experimental and CFD studies, the long flat surface between the front and rear of the body decouples the aerodynamic behaviour of these two regions and makes it more relevant to commercial vehicles. Two variants of the Windsor body were experimentally tested at the Loughborough University wind-tunnel [Varney2020a] and have been the subject of numerous CFD studies - that most recently have taken place within the 2nd, 3rd and 4th Automotive CFD Prediction Workshops [page2022]. The Windsor body was tested in numerous configurations but for this study we follow the 4th AutoCFD4 setup, where the vehicle has no wheels (which simplifies a common source of uncertainty) and the squareback rear shape is used, as the baseline geometry. As discussed later, this baseline geometry is then adapted according to 7 parameters to generate 355 different geometries.
@@
The model geometry is shown in 1. The reference frontal area is defined by the vehicle height and width and for the baseline geometry is . The reference length used for pitching moment is the wheelbase . The CAD geometry of the model has its origin on the ground plane, in the symmetry plane midway between the wheels (which are not included in this particular geometry) i.e . This is also the moment reference center for the calculating of the moments.
@@
All simulations were run using the Volcano ScaLES code by Volcano Platforms, which solves the explicit compressible Navier–Stokes equations using a nominally 4th order spatially accurate finite difference discretization with favorable Kinetic Energy and Entropy Consistency properties making it suitable for Large Eddy Simulations (LES) over a range of flow regimes. The viscous flux discretization utilizes a mix of 4th order and 2nd order discretizations with high spectral bandwidth thereby allowing further robustness for high Reynolds number flows without additional numerical dissipation via the inviscid flux numerical operator in turbulence resolving regions. Physics-based numerical sensors that are functions of local velocity gradients as well as pressure and density fluctuations allow for spatio-temporally localized use of limiters needed to capture flows with discontinuities such as shocks and steep density gradients. The Strong Stability Preserving (SSP) 3rd order Runge–Kutta scheme developed by Gottlieb & Shu [gottlieb1998total] is utilized for all the work presented in this paper.
Geometries are represented in the numerical formulation using an immersed boundary algorithm capable of enforcing inviscid no-penetration boundary conditions as well as viscous skin friction with appropriate Reynolds number asymptotic properties using an equilibrium wall-model. The wall-model that provides a shear-stress constraint at the wall uses solution information interpolated via probing at a fixed distance of into the fluid away from walls, where is the local grid spacing at the wall. Additional methods to damp generation of spurious numerical noise at walls as well as at coarse-fine interfaces are also utilized in the present work. All simulations presented in this work utilize the constant coefficient Vreman subgrid scale closure [vreman2004eddy] to model the subgrid-scale stresses.
@@
To follow the setup of the AutoCFD 3 and 4 workshops, the Windsor body is yawed 2.5 degrees and the Reynolds number is based on the body length and the freestream velocity. A nominal inlet velocity of is given upstream of the car and a sensor is placed at to measure local static pressure and velocity. For the validation of the method against experimental data for the baseline geometry, these are used to normalize the simulation outputs, however for the dataset itself, the reference inlet velocity is used to normalize all outputs (see SI for further details),. The remaining boundary conditions are shown in Figure 2.
@@
Only the ground plane has a no slip condition and hence has boundary layer growth; the top and side walls should be treated as a slip or “inviscid” wall.
@@
[…]
@@
| Label | Min. grid spacing | No. of cells | Wall time / CTU | |
@@
For the design case study, we choose to keep the mesh refinement regions fixed for all design cases. Using the fixed refinement topology, we consider a consistent grid refinement study for the Windsor body. An example of a grid in this family is depicted in Figure 3. The surface of the body is refined to the finest level with a large wake block immediately behind the body. The block is taller than necessary for the Windsor body but since we are choosing to keep the refinement regions fixed we define a region to accommodate all designs ride height. Table 1 shows results of a consistent grid refinement study using a sequence of grids for and . The columns of the table show the minimum grid spacing used, the number of computational cells in the grid, the wall-clock time per convective time unit333Nondimensional time is defined to be where is the length of the Windsor body, is the free-stream velocity, and is the time. (CTU) in minutes and the drag coefficient . An additional row is included for the experimental drag coefficient value. Based on the results, is chosen as a representative baseline case and the remainder of the experimental comparisons are carried out for the grid. This was to achieve a balance between accuracy, computational cost and reach a mesh size that is typically used by the automotive industry for external aerodynamics.
Velocity: Typical of bluff-bodies, the Windsor body shows a large recirculation region at the base. The data for the slice was taken with a freestream velocity of which is the same as our simulation but, in addition to this plane, tomographic PIV was taken for a 3D region behind the car. This data was recorded for a flow with a freestream velocity of . To make comparisons with the tomographic PIV data, we normalize the velocities by their respective freestream velocity. With that caveat, Figures 4 & 5 shows a comparison of the normalized time-averaged streamwise velocity for compared with the PIV measurements and the same solution at compared to the tomographic PIV data. The overall shape of the wake region is well captured as well as is its interaction with the floor boundary layer which can be seen in the velocity contours. See SI for a detailed discussion on the validation of the underlying CFD method.
@@
The baseline geometry is transferred into a parameterized CAD model to allow for automated sampling in the design space. The chosen sampling algorithm is a Halton sequence, which results in a quasi-random low discrepancy sequence with the advantages of being deterministic and extendable. The parameters and respective ranges are shown in Figure 6 and Table 2. The parameters were chosen to provide a suitable range of geometries that would exhibit different flow patterns e.g pressure vs geometry induced separation. The min/max values of each value were based upon engineering judgement and to avoid completely unrealistic shapes as well as avoid invalid shapes during geometry creation. The choice of 355 geometries were also based upon a mixture of computational budget as well as matching what may be possible within industry i.e. how many geometries would realistically be generated within an engineering company. Future work could be to expand this dataset further if required.
@@
The dataset is open-source under the CC-BY-SA license 444https://creativecommons.org/licenses/by-sa/4.0/) and available to download via Amazon S3. In order to download, no AWS account is required and the full details are provided in the SI, the dataset README 555https://caemldatasets.s3.us-east-1.amazonaws.com/windsor/dataset/README.txt and the website 666https://caemldatasets.org. Further mirroring sites are being explored and details will be on the website once available.
@@
The dataset follows the same structure as two concurrently developed datasets; AhmedML [ashton2024ahmed] and DrivAerML [ashton2024drivaer]. A summary of the dataset is provided in Table 3 (see SI for full details). The purpose of the dataset is to provide a rich dataset to support the development of a broad range of potential ML approaches. For this reason, we provide all possible outputs such as the full volume flow-field, the boundary surface, images of the flow-field as well as the time-averaged force and moment coefficients. All the outputs are provided in commonly used open-source formats (vtp, vtu, etc.) which allow for the the broadest compatibility and ease of use for developers.
@@
ML evaluation
We conducted an example ML evaluation using a Graph Neural Network (GNN) approach, based upon a modified version of MeshGraphNets [pfaff2021learning] (more details in the SI) to demonstrate how this dataset could be used to train a ML model to predict unseen cases. We assess two approaches; firstly directly predicting the Key Performance Indicators (KPIs); in this case the drag and lift coefficient values, for each geometry using the lower resolution STL surface mesh (89k nodes) as an input, and secondly predicting the lift and drag coefficient through integration of the surface wall-shear stress and pressure from the higher resolution VTP surface mesh (4.4M nodes). Using the first method we find that using a 60/20/20 split of train, validation and test data, it is possible to obtain a MSE of less than 0.00028 for the drag coefficient and a MSE less than 0.0175 for the lift coefficient. An example of the prediction accuracy is shown in Figure 8 for the training, validation and test data for the drag coefficient. Using x8 Nvidia L40s GPUs (Amazon EC2 g6e.48xlarge instances via Amazon Web Services), the training completes in approximately 2hrs and the inference time for each new predicted geometry is 0.15 seconds. Additional results and further details of the ML setup are provided in the SI. Please note that these ML evaluations are preliminary and purely serve to illustrate how this dataset can be used for ML evaluation. We hope other groups will use this dataset to do a more thorough evaluation of different ML methodologies.
@@ A.1 Large Eddy Simulations
Note that the dynamic viscosity, can be expressed as a function of Favre-filtered static temperature using an appropriate law (such as the Sutherland’s law). However, in the present work, this dynamic viscosity has be assumed to a spatio-temporal constant with a value determined to achieve the required Reynolds number. Static temperature, is a function of the local pressure and density in accordance to the ideal gas law using a gas constant, :
@@
where the unclosed terms on the right hand side represent interactions between subgrid-scale energy and the resolved scale energy. Since these terms are only relevant at high Mach numbers, the residual terms in the energy equation are modelled via just a turbulent Prandtl number and the Vreman’s eddy viscosity model used for the momentum equations. The positive definiteness of the SGS eddy viscosity ensures that the numerical model for the filtered energy equation is purely dissipative, i.e. it assumes that the energy is exclusively transferred from the large scales to the small scales of motion.
@@ B.1 Licensing terms
The dataset is provided with the Creative Commons CC-BY-SA v4.0 license777https://creativecommons.org/licenses/by-sa/4.0/deed.en. A full description of the license terms is provided under the following URL:
@@ B.3 Long-term hosting/maintenance plan
The data is hosted on Amazon S3 as it provides high durability, fast connectivity and is accessible without first requesting an account or credentials (only AWS CLI tools, described above, which are free to download and use). A dedicated website has been created for this dataset and the two associated datasets; AhmedML [ashton2024ahmed] and DrivAerML datasets [ashton2024drivaer] at https://caemldatasets.org to help further clarify where the data is hosted and to communicate any additional mirroring sites.
@@ B.7 Geometry variants
The geometry choice within the dataset results in a broad range of flow physics - that is partly illustrated in Figures 12 and 13 which show the large range of drag and lift coefficients across the dataset. Figures 10 and 11 take two examples from the dataset that shows the resulting flow-field (mean streamwise velocity) for a geometry producing a high drag coefficient and another one for a low drag coefficient.
@@
Figures 14, 15, 16, & 17 are provided to give an illustration of the range of flow features and post-processing that makes this a rich dataset for the development and testing of ML methods.
@@ Appendix C Additional Validation details
Pressure: In Figure 18 and Figure 19, we show experimental and computed surface pressure coefficient values, , for the cut planes and respectively. For the plane we observe stronger suction in the CFD compared to the experiment near where the flow transitions to turbulence, which is typical in other CFD simulations for this problem. Otherwise, there is good agreement for the upper and lower sections. For the cut, the CFD again shows a stronger suction peak this time in the A-pillar area. Small variations in the data suggest a longer averaging could be done but it would not be expected to change the agreement. Overall, the agreement is reasonably good and similar to other results seen in the AutoCFD3 results. A comparative view of the base pressure is shown in Figure 20 where similar pressure distributions are seen with a low pressure region on the leeward side of the vehicle associated with the recirculation region. The simulation results show a higher pressure than the experiment which is most evident on the side, however, the spatial distribution of the pressure and thus the shape of the contours match well.
@@
Velocity: Time-averaged streamlines colored by normalized mean streamwise velocity at are shown in Figure 21 compared with experimental data from the thesis by Varney [Varney2020]. The streamlines show the large recirculation region behind the vehicle and the correct flow topology. The data for the slice was taken with a freestream velocity of which is the same as our nominal freestream velocity but, in addition to this plane, tomographic PIV was taken for a 3D region behind the car. This data was recorded for a flow with a freestream velocity of . To make comparisons with the tomographic PIV data, we normalize their velocities by their freestream velocity. Figures 22 and 23 show progressive cuts through the wake going downstream. Starting nearest the car, the experiment shows a thicker shear layer but as we move downstream the predicted and experimental flow have essentially the same shear layer thickness. The overall shape of the wake region is well captured as well as its interaction with the floor boundary layer which can be seen in the velocity contours.
Appendix D ML evaluation
D.1 Method
We have conducted preliminary analysis on our dataset using a modified version of one of the state-of-the-art scientific machine learning (SciML) methods, MeshGraphNets [pfaff2021learning] on various tasks to illustrate the practicality of the dataset for ML evaluation. In both our method and MeshGraphNets, the node encoder and the edge encoder are 2-layer multilayer perceptrons (MLPs), and the processor consists of L message passing blocks, each containing two MLPs. There are a number of key differences between the approach taken in this work and that in the original paper of Pfaff et al.
The original justification and proof of MeshGraphNets was for transient prediction, which involves predicting dynamic quantities of the mesh at time t+1 given the current state and previous states. In contrast, our method focuses on directly predicting simulation results at the steady state or time-averaged state. We believe steady state or time-averaged prediction is more applicable in industrial contexts, where transient simulation data is often not preserved due to the large amount of data storage required.
This difference in task leads to distinct input node features. MeshGraphNet leverages dynamical features such as the instantaneous velocity or pressure to predict future states, but those features are not available in a steady state or time-averaged prediction. As our method predicts time-averaged results directly from meshes, the node features comprise node positions, node normals, and node defects in some use cases. Node positions represent the absolute node positions in the 3D space, as our method operates on 3D meshes. This is in contrast to the MeshGraphNets paper, which primarily focused on 2D CFD use cases. Our edge features, on the other hand, include the relative displacement vector (i.e., the difference between the positions of the two nodes forming the edge) and its norm, which closely resemble the edge features in MeshGraphNets.
Our model architecture has been adapted to address two distinct use cases: direct Key Performance Indictor (KPI) prediction and surface variable prediction. Among the two use cases, the architecture used in surface variable prediction more closely resembles that of MeshGraphNets, as both produce node level predictions. For KPI prediction, the models make graph level predictions, that is predicting one or a few values (e.g lift or drag coefficient) per graph rather than per node. To achieve this, there is a pooling layer being added between the processor and the decoder. The pooling layer converts node embeddings to graph embeddings which are then passed to the decoder.
D.2 Results
The entire WindsorML dataset is split into training (60%), validation (20%), and test (20%) sets. For each use case, the model is trained on the training set, and the checkpoint that had the best validation error was used to obtain the inference results on the test set.
Using the first method to directly predict the lift and drag coefficients we find that using a 60/20/20 split of train, validation and test data, it is possible to obtain a MSE of less than 0.00028 for the drag coefficient and a MSE less than 0.0175 for the lift coefficient. An example of the prediction accuracy is shown in Figure 25 & 26 for the training, validation and test data for the lift and drag coefficient. Using x8 Nvidia L40s GPUs (Amazon EC2 g6e.48xlarge instances via Amazon Web Services), the training completes in approximately 2hrs and the inference time for each new predicted geometry is 0.15 seconds. We ran this training for 1000 epochs using 5 message passing layers with a batch size of 4. The Adams optimizer was used with an initial learning rate of 0.003.
For the second method, where the lift and drag coefficient are obtained through the integration of the predicted surface pressure and wall-shear stress on the 4.4M node VTP surface mesh, we obtain predictions for the drag coefficient with a mean absolute error (MAE) of 0.03269 and a weighted mean absolute percentage error (WMAPE) of 0.1040. For the lift coefficient the mean absolute error (MAE) is 0.1854 and the weighted mean absolute percentage error (WMAPE) of 0.698 (as shown in Figures 27 & 28). Training time is approximately 60 hours on x8 NVidia L40s GPUs (Amazon EC2 g6e.48xlarge instances via Amazon Web Services) and the inference time is less than a minute on the same hardware. We ran this training with BF16 for 1200 epochs using 20 message passing layers with a batch size of 1. The Adams optimizer was used with an initial learning rate of 0.001.
It is clear from these results that the WindsorML dataset represents a challenging use-case for ML methods to predict, with noticeable difference in accuracy based upon the method used. Please note that these ML evaluations are preliminary and purely serve to illustrate how this dataset can be used for ML evaluation. We hope other groups will use this dataset to do a more thorough evaluation of different ML methodologies.
Appendix E Datasheet
E.1 Motivation